# Robust 6DoF Pose Estimation Against Depth Noise and a Comprehensive Evaluation on a Mobile Dataset

### v4

@@ 5 Experiments
depth-ADD DenseFusion(Wang et al., 2019) MegaPose-RGBD(Labbé et al., 2022) ES6D*(Mo et al., 2022) BundleSDF (Wen et al., 2023a) DTTDNet (Ours)
@@
Figure 6: Qualitative evaluation of different methods. To further validate our approach, we provide visual evidence of our model’s effectiveness in challenging occlusion scenarios and varying lighting conditions, where other models’ predictions fail but ours remain reliable. It should be noted that BundleSDF(Wen et al., 2023a) fails to reconstruct 3D objects in some scenes, resulting in the absence of annotations for such objects.
@@ 5.1 Evaluation Metrics
We evaluate baselines with the average distance metrics ADD and ADD-S according to previous protocols(Xiang et al., 2018; Feng et al., 2023). Suppose $R$ and $t$ are ground truth rotation and translation and $\tilde{R}$ and $\tilde{t}$ are the predicted counterparts. The ADD metric computes the mean of the pairwise distances between the 3D model points using ground truth pose $(R,t)$ and predicted pose $(\tilde{R},\tilde{t})$:
@@
Following previous protocols (Xiang et al., 2018; Wang et al., 2019; Liu et al., 2021, 2023; Labbé et al., 2022), a 3D pose estimation is deemed accurate if the average distance error falls below a predefined threshold. Two widely-used metrics are employed in our work, namely ADD/ADD-S AUC and ADD/ADD-S(1cm). For commonly used ADD/ADD-S AUC, we calculate the Area Under the Curve (AUC) of the success-threshold curve over different distance thresholds, where the threshold values are normalized between 0 and 1. On the other hand, ADD/ADD-S(1cm) is defined as the percentage of pose error smaller than the 1cm threshold.
@@
depth-ADD DenseFusion(Wang et al., 2019) MegaPose-RGBD(Labbé et al., 2022) ES6D(Mo et al., 2022) BundleSDF (Wen et al., 2023a) DTTDNet (Ours)
@@ 5.2 Experimental Results
In this section, we compare the performance of our method DTTDNet with four other 6DoF pose estimators, namely, BundleSDF(Wen et al., 2023a), MegaPose(Labbé et al., 2022), ES6D(Mo et al., 2022), DenseFusion(Wang et al., 2019). While all four methods leverage the benefits of multimodal data from both RGB and depth sources, they differ in the extent to which they emphasize the depth data processing module. Quantitative experimental results are shown in Table 2. Qualitative examples are shown in Fig. 6.
BundleSDF(Wen et al., 2023a) learns multi-view consistent shape and appearance of a 3D object using an object-centric neural signed distance field, leverages frame poses captured in-flight. However, the approach struggles with 3D object reconstruction when faced with large distances, low resolution, or insufficient frame sequence per viewpoint. BundleSDF(Wen et al., 2023a) achieves an ADD AUC of 46.86 and an ADD-S AUC of 55.74. This method failed to reconstruct 8 out of 26 object-scene combinations in DTTD-Mobile. As shown in 6, the failure in 3D object reconstruction results in the absence of object pose tracking in the corresponding scenes. The BundleSDF, as shown in Figure 1, excludes objects in scenes lacking associated pose tracking.
MegaPose(Labbé et al., 2022) employs a coarse-to-fine process for pose estimation. The initial ”coarse” module leverages both RGB and depth data to identify the most probable pose hypothesis. Subsequently, a more precise pose inference is achieved through the ”render-and-compare” technique. Disregarding the noise in the depth data can also impair the effectiveness of their coarse module, consequently leading to failure in their refinement process. Even with the assistance of a refiner, MegaPose-RGBD(Labbé et al., 2022) only manages to attain an ADD AUC of 49.02 and an ADD-S AUC of 62.44. Its damage and susceptibility to depth noise falls somewhere between DenseFusion(Wang et al., 2019) and ES6D(Mo et al., 2022).
DenseFusion(Wang et al., 2019) treats both modalities equally and lacks a specific design for the depth module, whereas ES6D(Mo et al., 2022) heavily relies on depth data during training, using grouped primitives to prevent point-pair mismatch. However, due to potential interpolation errors in the depth data, this additional supervision can introduce erroneous signals to the estimator, resulting in inferior performance compared to DenseFusion(Wang et al., 2019). DenseFusion(Wang et al., 2019) achieves 69.67 ADD AUC and 85.88 ADD-S AUC, whereas ES6D(Mo et al., 2022) only achieves 13.25 ADD AUC and 37.38 ADD-S AUC.
@@ 5.3 Ablation Studies
Evaluation of DTTDNet on other datasets. To show that our proposed pose estimator could also be generalized to other domains, we evaluate our method on the YCB-Video dataset(Xiang et al., 2018), which outperforms MegaPose(Labbé et al., 2022) and ES6D(Mo et al., 2022) by 6.87 and 7.90 points on ADD(S), respectively, when performing fair comparison (i.e., without any data pre-cleaning and iterative refinement).
We share 3 overlapping objects between the YCB video dataset and our DTTD-Mobile dataset, and demonstrate the models’ performance on these objects to examine the performance drop of each baseline when shifting the datasets, as shown in Table 3. DTTDNet achieves the highest performance on both datasets, and shows less performance drop when occurring iPhone LiDAR noise. Detailed per-object evaluation is appended in the supplementary materials.
@@
Effect of Depth Feature Filtering module. Table 4 illustrates the improvement in ADD AUC metrics achieved by our method when integrating geometric feature filtering (GFF) module. To provide a detailed insight into the impact of the GFF module, we conducted principal component analysis (PCA) on both the initial geometric tokens encoded by the PointNet and the filtered version after applying the GFF module, i.e., projected the embedding to a 1-D array with its dominant factor. We visualize the geometric embedding both before and after the application of the GFF module by generating histograms of the dimensionally reduced geometric tokens, as shown in Fig. 8. The distribution of these tokens, as shown in the right subplot, becomes more balanced and uniform after learning-based filtering through the GFF module. The enhanced ADD AUC performance can be attributed to the balanced distribution achieved through the use of the depth robustifying module.
@@
Effect of Layer Number Variation in Fusion Stages. Table 6 display the variations brought about by increasing the number of layers at different fusion stages. Overall, adding layer number increases the model’s performance in terms of ADD AUC. As we proportionally increase the total number of layers in the modality fusion or point-wise fusion, we witness a sustained improvement in model performance.
@@ 6 Conclusion
Appendix A Related Work
@@
6DoF Pose Estimation Algorithms. The majority of data-driven approaches for object pose estimation revolve around utilizing either RGB images (Labbe et al., 2020; Peng et al., 2019; Xiang et al., 2018; Zakharov et al., 2019) or RGBD images (He et al., 2021c, 2020; Wang et al., 2019; Mo et al., 2022; Jiang et al., 2022) as their input source. RGBD remains mainstream in industrial environments requiring higher precision. However, due to the high cost of accurate depth sensors, finding more robust solutions compatible with inexpensive and widely used sensors is a problem we hope to address.
@@
Methods(He et al., 2021c, 2020; Wang et al., 2019; Mo et al., 2022; Jiang et al., 2022) that relied on depth maps advocated for the modality fusion of depth and RGB data to enhance their inference capabilities. To effectively fuse multi-modalities, Wang et al. (Wang et al., 2019) introduced a network architecture capable of extracting and integrating dense feature embedding from both RGB and depth sources. Due to its simplicity, this method achieved high efficiency in predicting object poses. In more recent works (He et al., 2020, 2021c, 2022), performance improvements were achieved through more sophisticated network architectures. For instance, He et al. (He et al., 2021c) proposed an enhanced bidirectional fusion network for key-point matching, resulting in high accuracy on benchmarks such as YCB-Video (Xiang et al., 2018) and LINEMOD (Hinterstoisser et al., 2013). However, these methods exhibited reduced efficiency due to the complex hybrid network structures and processing stages. Addressing symmetric objects, Mo et al. (Mo et al., 2022) proposed a symmetry-invariant pose distance metric to mitigate issues related to local minima. On the other hand, Jiang et al. (Jiang et al., 2022) proposed an L1-regularization loss named abc loss, which enhanced pose estimation accuracy for non-symmetric objects.
@@
Besides the RGBD approach, studies following the RGB-only approach often rely on incorporating additional prior information and inductive biases during the inference process. These requirements impose additional constraints on the application of 3D object tracking on mobile devices. Their inference process can involve utilizing more viewpoints for similarity matching(Liu et al., 2023; Labbé et al., 2022) or geometry reconstruction(Sun et al., 2022; Wen et al., 2023b), employing rendering techniques(Labbé et al., 2022; Nguyen et al., 2022; Cai et al., 2022) based on precise 3D model or leveraging an additional database for viewpoint encoding retrieval(Cai et al., 2022). During the training phase, these approaches typically draw upon more extensive datasets, such as synthetic datasets, to facilitate effective generalization within open-set scenarios. However, when confronted with a limited set of data samples, their performance does not surpass that of closed-set algorithms in cases where there is a surplus of prior information available and depth map loss.
@@
3D Object Tracking Datasets Existing object pose estimation algorithms are predominantly tested on a limited set of real-world 3D object tracking datasets (Hinterstoisser et al., 2013; Marion et al., 2018; Xiang et al., 2018; Liu et al., 2020; Hodaň et al., 2017; Liu et al., 2021; Feng et al., 2023; Calli et al., 2015a, b, 2017), which often employ depth-from-stereo sensors or time-of-flight (ToF) sensors for data collection. Datasets like YCB-Video(Xiang et al., 2018), LINEMOD(Hinterstoisser et al., 2013), StereoOBJ-1M(Liu et al., 2021), and TOD(Liu et al., 2020) utilize depth-from-stereo sensors, while TLess(Hodaň et al., 2017) and DTTD (Feng et al., 2023) deploy ToF sensors, specifically the Microsoft Azure Kinect, to capture meter-scale RGBD data. However, the use of cameras with depth-from-stereo sensors may not be an optimal platform for deploying AR software, because stereo sensors may degrade rapidly in longer-distance (Haggag et al., 2013) and may encounter issues with holes in the depth map when stereo matching fails. In our pursuit of addressing the limitations of existing datasets and ensuring a more realistic dataset captured with mobile devices, we opt to collect RGBD data using the iPhone 14 Pro.
iPhone-based Datasets for 3D Applications. Several datasets utilize the iPhone as their data collection device for 3D applications, such as ARKitScenes(Baruch et al., 2021), MobileBrick(Li et al., 2023a), ARKitTrack(Zhao et al., 2023), and RGBD Dataset(He et al., 2021b). These datasets were constructed to target applications from 3D indoor scene reconstruction, 3D ground-truth annotation, depth-map pairing from different sensors, to RGBD tracking in both static and dynamic scenes. However, most of these datasets did not specifically target the task of 6DoF object pose estimation. Our dataset provides a distinct focus on this task, offering per-pixel segmentation and pose labels. This enables researchers to delve into the 3D localization tasks of objects with a dataset specifically designed for this purpose. The most relevant work is from OnePose(Sun et al., 2022), which is an RGBD 3D dataset collected by iPhone. However, their dataset did not provide 3D models for close-set settings, and they utilized automatic localization provided by ARKit for pose annotation, which involved non-trivial error for high-accuracy 6DoF pose estimation. On the other hand, we achieve higher localization accuracy with the OptiTrack professional motion capture system to track the iPhone camera’s real-time positions as it moves in 3D.
Appendix B More Dataset Description
B.1 Data Acquisition
@@
To track the iPhone’s 6DoF movement, we did not use the iPhone’s own world tracking SDK. Instead, we follow the same procedure as in (Feng et al., 2023) and use the professional OptiTrack motion capture system for higher accuracy. For label generation, we also use the open-sourced data annotation pipeline provided by (Feng et al., 2023) to annotate and refine ground-truth poses for objects in the scenes along with per-pixel semantic segmentation. Some visualizations of data samples are illustrated in Fig. 9. Notice that the scenes cover various real-world occlusion and lighting conditions with high-quality annotations. Following previous dataset protocols (Feng et al., 2023; Xiang et al., 2018), we also provide synthetic data for scene augmentations used for training.
@@
B.2 Train/Test Split
DTTD-Mobile offers a suggested train/test partition as follows. The training set contains 8622 keyframes extracted from 88 video sequences, while the testing set contains 1239 keyframes from 12 video sequences. To ensure a representative distribution of scenes with occluded objects and varying lighting conditions, we randomly allocate them across both the training and testing sets. Furthermore, for training purposes of scene augmentations, we provide 20,000 synthetic images by randomly placing objects in scenes using the data synthesizer provided in (Feng et al., 2023).
Appendix C More Implementation Details
C.1 Details on RGBD Feature Fusion
@@
Modality Fusion. The objective of this module is to combine geometric embedding $g$ and RGB embedding $c$ produced by single-modal encoders in a cross-modal fashion. Drawing inspiration from ViLP (Kim et al., 2021), both types of embedding are linearly transformed into a token sequence ($\in\mathbb{R}^{N\times d_{emb}}$). Before entering the modality fusion module $E_{1}$, these features are combined along the sequence length direction, i.e., all feature embedding is concentrated into a single combined sequence, where the dimension remains $d_{emb}$, and the sequence length becomes twice the original length.
@@
where the operation symbol $\oplus$” denotes concentrating along the row direction. It is then reshaped into the sequence $f_{1}^{\prime}$ with the length of N and dimension of $2d_{f_{1}}$ in order to adapt the point-wise transformer encoder in the next fusion stage. This step enables the model’s attention mechanism to effectively perform cross-modal fusion tasks.
Point-Wise Fusion. The goal of this stage is to enhance the integration of information among various points. The primary advantage of our method over the previous work(He et al., 2021c) is that our model can calculate similarity scores not only with the nearest point but also with all other points, allowing for more comprehensive interactions. In order to enable the point-wise fusion to effectively capture the similarities between different points, we merge the original RGB token sequence $c$ and the geometric token sequence $g$ together with the output embedding sequence $m^{\prime}$ from the modality fusion module along the feature dimension direction. The combined sequence input $\left[c^{T}\oplus g^{T}\oplus(f_{1}^{\prime})^{T}\right]^{T}\in\mathbb{R}^{(2d_{emb}+2d_{f_{1}})\times N}$ is then fed into the point-wise transformer encoder $E_{2}$ to acquire the final fusion:
@@
C.2 Hyper Parameters
Training Strategies. For our DTTDNet, learning rate warm-up schedule is used to ensure that our transformer-based model can overcome local minima in early stage and be more effectively trained. By empirical evaluation, in the first epoch, the learning rate $lr$ linearly increases from 00 to $1e{-5}$. In the subsequent epochs, it is decreased using a cosine scheduler to the end learning rate $min\_lr=1e{-6}$. Additionally, following the approach of DenseFusion(Wang et al., 2019), we also decay our learning rate by a certain ratio when the average error is below a certain threshold during the training process. Detailed code and parameters will be publicly available in our code repository. Moreover, we set the importance factor $\lambda$ of CDL to $0.3$ and the initial balancing weight $w$ to $0.015$ by empirical testing.
Appendix D More Experimental Details
D.1 Baseline Implementation Details
For all the baseline methods(Wang et al., 2019; Labbé et al., 2022; Mo et al., 2022; Wen et al., 2023a) that we adopted, we did not integrate any additional iterative refinement processes (e.g., ICP(Besl & McKay, 1992)) for fair comparison.
ES6D(Mo et al., 2022). We preprocessed the training datasets including both DTTD-Mobile and YCB video according to the original paper and official codebase of ES6D(Mo et al., 2022), including removing some high-noise data, normalizing 3D translation, averaging the xyz map, and filtering out outliers in the point cloud. For the test set, to ensure fair comparison with other baselines, we did not adopt certain noise reduction methods that require prior knowledge based on ground truth data, nor did we exclude some high-noise data samples. We ensured that all test samples were retained and had errors calculated as those in other baselines.
BundleSDF(Wen et al., 2023a). The object-centric camera pose coordinates outputted by BundleSDF(Wen et al., 2023a) are based on its own embedded geometric reconstruction. In order to compute metrics in the same coordinate system and with the same CAD models as other baselines, we aligned the camera trajectory computed for each scene-object combination with the ground truth camera pose through trajectory-wise alignment based on the Umeyama algorithm (Umeyama, 1991).
D.2 More Ablation Studies and Analysis
@@
Attention Map Visualization. To visualize what our fusion module learns during the training process, we draw on previous studies (Trockman & Kolter, 2023; Guo et al., 2023) and represent our attention map as $a_{i,j}$ described in section C.1. Taking two objects (itoen_green_tea and black_marker) as examples, Fig. 10 displays the attention maps produced by different attention heads in the two fusion stages. We showcase the attention maps generated by the modality fusion and point-wise fusion at their respective final layers. The modality fusion part reveals distinct quadrant-like patterns, reflecting differences in how the two modalities fuse. The lower-left and upper-right quadrants offer insights into the degree of RGB and geometric feature fusion. The point-wise fusion part exhibits a striped pattern and shows that it attends to the significance of specific tokens during training.
D.3 More Results on YCB video Dataset
Due to the lower depth noise in the YCB video dataset (YCB’s average depth-ADD is 0.009, while DTTD-Mobile’s average depth-ADD is 0.239), we adopted a simplified DTTDNet model structure, omitting the GFF module, to seek faster training convergence. Additionally, the reference point set used for computing Chamfer distance loss was directly extracted from the depth map. Furthermore, regarding hyper-parameters, we chose 0 layers for modality fusion and 6 self-attention layers for point-wise fusion.
When examining the performance of each baseline on the YCB video dataset, as shown in 2. DTTDNet achieves the highest average performance, with an ADD-S AUC of 94.19 and an ADD-S (2cm) of 96.14. DenseFusion (Wang et al., 2019) is the second best, with an ADD-S AUC of 91.20 and an ADD-S (2cm) of 95.30. Although BundleSDF (Wen et al., 2023a) shows strong performance across many object classes, it struggles with pose estimation for some objects, primarily due to its inability to reconstruct 3D models in the presence of occlusions. Its ADD-S AUC is 86.31, and its ADD-S (2cm) is 87.64. In contrast, MegaPose-RGBD (Labbé et al., 2022) and ES6D (Mo et al., 2022) lag behind in performance, particularly in the number of object classes in which they perform the best, with ADD-S AUC scores of 82.64 and 78.14, and ADD-S (2cm) scores of 84.81 and 75.35, respectively.
@@ D.3 More Results on YCB video Dataset
Table 7: Comparison with diverse 6DoF pose estimation baselines on YCB video dataset. We evaluate the results as the prior works(Wang et al., 2019) using ADD-S AUC and ADD-S (2cm) on all 21 objects, higher is better. Note that the left-most column indicates the per-object depth-ADD error. Objects with bold names are symmetric.
depth-ADD DenseFusion(Wang et al., 2019) MegaPose-RGBD(Labbé et al., 2022) ES6D(Mo et al., 2022) BundleSDF (Wen et al., 2023a) DTTDNet (Ours)

### v5

@@ 5 Experiments
depth-ADD DenseFusion [51] MegaPose-RGBD [31] ES6D [39] BundleSDF [53] DTTDNet (Ours)
@@
Figure 6: Qualitative evaluation of different methods. To further validate our approach, we provide visual evidence of our model’s effectiveness in challenging occlusion scenarios and varying lighting conditions, where other models’ predictions fail but ours remain reliable. It should be noted that BundleSDF [53] fails to reconstruct 3D objects in some scenes, resulting in the absence of annotations for such objects.
@@ 5.1 Evaluation Metrics
We evaluate baselines with the average distance metrics ADD and ADD-S according to previous protocols [56, 12]. Suppose $R$ and $t$ are ground truth rotation and translation and $\tilde{R}$ and $\tilde{t}$ are the predicted counterparts. The ADD metric computes the mean of the pairwise distances between the 3D model points using ground truth pose $(R,t)$ and predicted pose $(\tilde{R},\tilde{t})$:
@@
Following previous protocols [56, 51, 36, 37, 31], a 3D pose estimation is deemed accurate if the average distance error falls below a predefined threshold. Two widely-used metrics are employed in our work, namely ADD/ADD-S AUC and ADD/ADD-S(1cm). For commonly used ADD/ADD-S AUC, we calculate the Area Under the Curve (AUC) of the success-threshold curve over different distance thresholds, where the threshold values are normalized between 0 and 1. On the other hand, ADD/ADD-S(1cm) is defined as the percentage of pose error smaller than the 1cm threshold.
@@
depth-ADD DenseFusion [51] MegaPose-RGBD [31] ES6D [39] BundleSDF [53] DTTDNet (Ours)
@@ 5.2 Experimental Results
In this section, we compare the performance of our method DTTDNet with four other 6DoF pose estimators, namely, BundleSDF [53], MegaPose [31], ES6D [39], DenseFusion [51]. While all four methods leverage the benefits of multimodal data from both RGB and depth sources, they differ in the extent to which they emphasize the depth data processing module. Quantitative experimental results are shown in Table 2. Qualitative examples are shown in Fig. 6.
BundleSDF [53] learns multi-view consistent shape and appearance of a 3D object using an object-centric neural signed distance field, leverages frame poses captured in-flight. However, the approach struggles with 3D object reconstruction when faced with large distances, low resolution, or insufficient frame sequence per viewpoint. BundleSDF [53] achieves an ADD AUC of 46.86 and an ADD-S AUC of 55.74. This method failed to reconstruct 8 out of 26 object-scene combinations in DTTD-Mobile. As shown in 6, the failure in 3D object reconstruction results in the absence of object pose tracking in the corresponding scenes. The BundleSDF, as shown in Figure 1, excludes objects in scenes lacking associated pose tracking.
MegaPose [31] employs a coarse-to-fine process for pose estimation. The initial "coarse" module leverages both RGB and depth data to identify the most probable pose hypothesis. Subsequently, a more precise pose inference is achieved through the "render-and-compare" technique. Disregarding the noise in the depth data can also impair the effectiveness of their coarse module, consequently leading to failure in their refinement process. Even with the assistance of a refiner, MegaPose-RGBD [31] only manages to attain an ADD AUC of 49.02 and an ADD-S AUC of 62.44. Its damage and susceptibility to depth noise falls somewhere between DenseFusion [51] and ES6D [39].
DenseFusion [51] treats both modalities equally and lacks a specific design for the depth module, whereas ES6D [39] heavily relies on depth data during training, using grouped primitives to prevent point-pair mismatch. However, due to potential interpolation errors in the depth data, this additional supervision can introduce erroneous signals to the estimator, resulting in inferior performance compared to DenseFusion [51]. DenseFusion [51] achieves 69.67 ADD AUC and 85.88 ADD-S AUC, whereas ES6D [39] only achieves 13.25 ADD AUC and 37.38 ADD-S AUC.
@@
To assess the real-time applicability of DTTDNet, we benchmarked its inference speed on a single NVIDIA RTX A6000 GPU. Our model achieves an average inference time of 0.0172 seconds per object and 0.0378 seconds per frame, corresponding to 58.01 objects per second and 26.43 FPS. These results demonstrate that DTTDNet is well-suited for real-time 6DoF pose estimation, achieving efficient performance without sacrificing robustness or accuracy.
@@ 5.3 Ablation Studies
Evaluation of DTTDNet on other datasets. To show that our proposed pose estimator could also be generalized to other domains, we evaluate our method on the YCB-Video dataset [56], which outperforms MegaPose [31] and ES6D [39] by 6.87 and 7.90 points on ADD(S), respectively, when performing fair comparison (i.e., without any data pre-cleaning and iterative refinement).
We share 3 overlapping objects between the YCB video dataset and our DTTD-Mobile dataset, and demonstrate the models’ performance on these objects to examine performance drop of each baseline when shifting the datasets, as shown in Table 3. DTTDNet achieves the highest performance on both datasets, and shows less performance drop when occurring iPhone LiDAR noise. Detailed per-object evaluation is appended in the appendix.
@@
Effect of Depth Feature Filtering module. Table 4 illustrates the improvement in ADD AUC metrics achieved by our method when integrating geometric feature filtering module. To provide a detailed insight into the impact of the GFF module, we conducted principal component analysis (PCA) on both the initial geometric tokens encoded by the PointNet and the filtered version after applying the GFF module, i.e., projected the embedding to a 1-D array with its dominant factor. We visualize the geometric embedding both before and after the application of the GFF module by generating histograms of the dimensionally reduced geometric tokens, as shown in Fig. 8. The distribution of these tokens, as shown in the right subplot, becomes more balanced and uniform after learning-based filtering through the GFF module. The enhanced ADD AUC performance can be attributed to the balanced distribution achieved through the use of the depth robustifying module.
@@
Effect of Layer Number Variation in Fusion Stages. Table 6 display the variations brought about by increasing the number of layers at different fusion stages. Overall, adding layer number increases the model’s performance in terms of ADD AUC. As we proportionally increase the total number of layers in the modality fusion or point-wise fusion, sustained improvement is observed.
@@ 6 Conclusion
7 Acknowledgment
This work is supported by the FHL Vive Center at the University of California, Berkeley. We gratefully acknowledge our other team members: Tianjian Xu and Kathy Zhuang for their support with the iPhone capturing app and BundleSDF evaluation on our dataset; Xiang Zhang for his assistance in the further robotic dataset extension [^4] DTTD-Fanuc is an annotated dataset specifically designed for robotic grasping and manipulation with cheaper and low-resolution sensors. It is publicly available in our DTTDNet GitHub repository.; Weiyu Feng and Chenfeng Xu for their valuable discussions and academic insights during the development of this project.
Supplementary Material
8 Related Work
@@
6DoF Pose Estimation Algorithms. The majority of data-driven approaches for object pose estimation revolve around utilizing either RGB images [30, 42, 56, 58] or RGBD images [20, 19, 51, 39, 26] as their input source. RGBD remains mainstream in industrial environments requiring higher precision. However, due to the high cost of accurate depth sensors, finding more robust solutions compatible with inexpensive and widely used sensors is a problem we aim to address.
@@
Methods [20, 19, 51, 39, 26] that relied on depth maps advocated for the modality fusion of depth and RGB data to enhance inference capabilities. To effectively fuse multi-modalities, Wang et al. [51] introduced a network architecture capable of extracting and integrating dense feature embedding from both RGB and depth sources. Due to its simplicity, this method achieved high efficiency in predicting object poses. In more recent works [19, 20, 21], performance improvements were achieved through more sophisticated network architectures. For instance, He et al. [20] proposed an enhanced bidirectional fusion network for key-point matching, resulting in high accuracy on benchmarks such as YCB-Video [56] and LINEMOD [23]. However, these methods exhibited reduced efficiency due to the complex hybrid network structures and processing stages. Addressing symmetric objects, Mo et al. [39] proposed a symmetry-invariant pose distance metric to mitigate issues related to local minima. On the other hand, Jiang et al. [26] proposed an L1-regularization loss named abc loss, which enhanced pose estimation accuracy for non-symmetric objects.
@@
Besides the RGBD approach, studies following the RGB-only approach often rely on incorporating additional prior information and inductive biases during the inference process. These requirements impose additional constraints on the application of 3D object tracking on mobile devices. Their inference process can involve utilizing more viewpoints for similarity matching [37, 31] or geometry reconstruction [47, 54], employing rendering techniques [31, 40, 5] based on precise 3D model or leveraging an additional database for viewpoint encoding retrieval [5]. During the training phase, these approaches typically draw upon more extensive datasets, such as synthetic datasets, to facilitate effective generalization within open-set scenarios. However, when confronted with a limited set of data samples, their performance does not surpass that of closed-set algorithms in cases where there is a surplus of prior information available and depth map loss.
@@
3D Object Tracking Datasets Existing object pose estimation algorithms are predominantly tested on a limited set of real-world 3D object tracking datasets [23, 38, 56, 35, 24, 36, 12, 6, 7, 8], which often employ depth-from-stereo sensors or time-of-flight (ToF) sensors for data collection. Datasets like YCB-Video [56], LINEMOD [23], StereoOBJ-1M [36], and TOD [35] utilize depth-from-stereo sensors, while TLess [24] and DTTD [12] deploy ToF sensors, specifically the Microsoft Azure Kinect, to capture meter-scale RGBD data. However, the use of cameras with depth-from-stereo sensors may not be an optimal platform for deploying AR software, because stereo sensors may degrade rapidly at longer distances [16] and may encounter issues with holes in the depth map when stereo matching fails. In our pursuit of addressing the limitations of existing datasets and ensuring a more realistic dataset captured with mobile devices, we opt to collect RGBD data using the iPhone 14 Pro.
iPhone-based Datasets for 3D Applications. Several datasets utilize the iPhone as their data collection device for 3D applications, such as ARKitScenes [2], MobileBrick [33], ARKitTrack [60], and RGBD Dataset [18]. These datasets were constructed to target applications from 3D indoor scene reconstruction, 3D ground-truth annotation, depth-map pairing from different sensors, to RGBD tracking in both static and dynamic scenes. However, most of these datasets did not specifically target the task of 6DoF object pose estimation. Our dataset provides a distinct focus on this task, offering per-pixel segmentation and pose labels. This enables researchers to delve into the 3D localization tasks of objects with a dataset specifically designed for this purpose. The most relevant work is from OnePose [47], which is an RGBD 3D dataset collected by iPhone. However, their dataset did not provide 3D models for close-set settings, and they utilized automatic localization provided by ARKit for pose annotation, which involved non-trivial error for high-accuracy 6DoF pose estimation. On the other hand, we achieve higher localization accuracy with the OptiTrack professional motion capture system to track the iPhone camera’s real-time positions as it moves in 3D.
9 More Dataset Description
9.1 Data Acquisition
@@
To track the iPhone’s 6DoF movement, we did not use the iPhone’s own world tracking SDK. Instead, we follow the same procedure as in [12] and use the professional OptiTrack motion capture system for higher accuracy. For label generation, we also use the open-sourced data annotation pipeline provided by [12] to annotate and refine ground-truth poses for objects in the scenes along with per-pixel semantic segmentation. Some visualizations of data samples are illustrated in Fig. 9. Notice that the scenes cover various real-world occlusion and lighting conditions with high-quality annotations. Following previous dataset protocols [12, 56], we also provide synthetic data for scene augmentations used for training.
@@
9.2 Train/Test Split
DTTD-Mobile offers a suggested train/test partition as follows. The training set contains 8622 keyframes extracted from 88 video sequences, while the testing set contains 1239 keyframes from 12 video sequences. To ensure a representative distribution of scenes with occluded objects and varying lighting conditions, we randomly allocate them across both the training and testing sets. Furthermore, for training purposes of scene augmentations, we provide 20,000 synthetic images by randomly placing objects in scenes using the data synthesizer provided in [12].
10 More Implementation Details
10.1 Details on RGBD Feature Fusion
@@
Modality Fusion. The objective of this module is to combine geometric embedding $g$ and RGB embedding $c$ produced by single-modal encoders in a cross-modal fashion. Drawing inspiration from ViLP [28], both types of embedding are linearly transformed into a token sequence ($\in\mathbb{R}^{N\times d_{emb}}$). Before entering the modality fusion module $E_{1}$, these features are combined along the sequence length direction, i.e., all feature embedding is concentrated into a single combined sequence, where the dimension remains $d_{emb}$, and the sequence length becomes twice the original length.
@@
where the operation symbol $\oplus$" denotes concentrating along the row direction. It is then reshaped into the sequence $f_{1}^{\prime}$ with the length of N and dimension of $2d_{f_{1}}$ in order to adapt the point-wise transformer encoder in the next fusion stage. This step enables the model’s attention mechanism to effectively perform cross-modal fusion tasks.
Point-Wise Fusion. The goal of this stage is to enhance the integration of information among various points. The primary advantage of our method over the previous work [20] is that our model can calculate similarity scores not only with the nearest point but also with all other points, allowing for more comprehensive interactions. In order to enable the point-wise fusion to effectively capture the similarities between different points, we merge the original RGB token sequence $c$ and the geometric token sequence $g$ together with the output embedding sequence $m^{\prime}$ from the modality fusion module along the feature dimension direction. The combined sequence input $\left[c^{T}\oplus g^{T}\oplus(f_{1}^{\prime})^{T}\right]^{T}\in\mathbb{R}^{(2d_{emb}+2d_{f_{1}})\times N}$ is then fed into the point-wise transformer encoder $E_{2}$ to acquire the final fusion:
@@
Attention Map Visualization. To visualize what our fusion module learns during the training process, we draw on previous studies [48, 15] and represent our attention map as $a_{i,j}$ described in section 10.1. Taking two objects (itoen_green_tea and black_marker) as examples, Fig. 10 displays the attention maps produced by different attention heads in the two fusion stages. We showcase the attention maps generated by the modality fusion and point-wise fusion at their respective final layers. The modality fusion part reveals distinct quadrant-like patterns, reflecting differences in how the two modalities fuse. The lower-left and upper-right quadrants offer insights into the degree of RGB and geometric feature fusion. The point-wise fusion part exhibits a striped pattern and shows that it attends to the significance of specific tokens during training.
@@
Table 7: Comparison with diverse 6DoF pose estimation baselines on YCB video dataset. We evaluate the results as the prior works [51] using ADD-S AUC and ADD-S (2cm) on all 21 objects, higher is better. Note that the left-most column indicates the per-object depth-ADD error. Objects with names in bold are symmetric.
depth-ADD DenseFusion [51] MegaPose-RGBD [31] ES6D [39] BundleSDF [53] DTTDNet (Ours)